OpenStudy (anonymous):

Solve please! \[\left| -2-x \right| =2\]

jimthompson5910 (jim_thompson5910):

\[\Large \left|-2-x\right|=2\] \[\Large -2-x=2 \ \text{or} \ -2-x=-2\] \[\Large -x=2+2 \ \text{or} \ -x=-2+2\] \[\Large -x=4 \ \text{or} \ -x=0\] \[\Large x=-4 \ \text{or} \ x=0\]